# A Secure Estimator with Gaussian Bernoulli Mixture Model

### v1

@@ VI-B Gaussian-Bernoulli Secure (GBS) estimator
$\arg\min_{\hat{p}_{0:N}}~{}\mathcal{W}(\hat{x}_{0:N}^{*},\hat{p}_{0:N})$
$=$ $\biggl{\{}\sum_{i=0}^{N}[(1-\hat{p}_{i})\|y_{i}-C\hat{x}_{i}^{*}\|^{2}_{R^{-1}}+\hat{p}_{i}\alpha]\biggr{\}}$ (22)
@@
We can update $\hat{p}_{k}$ by solving the following optimization problem:
@@
As described in Algorithm 3, we iteratively calculate equation (21) and (VI-B), and improve the optimality through equation (26).
Input: the system parameters $\{A,C,Q,R,\alpha,\tau\}$
Input: the initial state estimate $(\bar{x}_{0},\bar{P}_{0})$ and the full observation set $y_{0:N}$
Output: $\hat{x}_{0:N},\hat{p}_{0:N}$
1 Initialize $\hat{x}_{0:N}^{0},\hat{p}_{0:N}^{0}$ using Algorithm 2 ;
2 Set the count $k=0$ and ${\mathcal{O}^{0}}=\{i|\hat{p}_{i}^{0}=0\}$;
3 repeat
4       $k\leftarrow k+1$;
       // $\arg\min_{\hat{p}_{0:N}}~{}\mathcal{L}(\hat{x}_{0:N}^{*},\hat{p}_{0:N})$
5       for $i=0$ to $N$  do
6             $\hat{p}_{i}^{k}=\textbf{bool}(\|y_{i}-C\hat{x}_{i}^{*}\|^{2}_{R^{-1}}>\alpha)$;
7
8      update $\mathcal{O}^{k}=\{i|\hat{p}_{i}^{k}=0\}$;
       // check suboptimal points
9       if $\mathcal{O}^{k}=\mathcal{O}^{k-1}$ then
10             for $i=0$ to $N$ do
11                   $\hat{p}_{i}^{*}=\textbf{bool}((\min\mathcal{L}_{\mathcal{O}_{i}^{+}}-\min\mathcal{L}_{\mathcal{O}_{i}^{-}})>\alpha)$;
12                   if $\hat{p}_{i}^{*}\neq\hat{p}_{i}^{k}$ then
13                         $\hat{p}_{i}^{k}=\neg\hat{p}_{i}^{k}$ and update $\mathcal{O}^{k}$ ;
14                         break;
15
16
17
      // $\arg\min_{\hat{x}_{0:N}}~{}\mathcal{L}(\hat{x}_{0:N},\hat{p}_{0:N}^{*})$
18       if $\mathcal{O}^{k}=\mathcal{O}^{k-1}$ then
19             break;
20
21      else
22             update optimal estimates $\hat{x}_{0:N}$;
23
24
25until ;
// judge status of observations
26 if $|\mathcal{O}^{k}|<(N+1-\tau)$ then
27       Print: info(”attack alarm in this sensor!”);
28
@@ 27       Print: info(”attack alarm in this sensor!”);
VI-C Convergence analysis
Here, we prove the convergence of our GBS estimator algorithm.
First, we demonstrate the existence of the solution to our target MIP problem presented in (7). As a $0-1$ sequence $\hat{p}_{0:N}\in\{0,1\}^{(N+1)}$, there are a total of $2^{(N+1)}$ cases. For $\{0,1\}^{(N+1)}\leftrightarrow 2^{\{0,\cdots,N\}}$, we can define a mapping from the indicator set to the index set that
@@ VI-C Convergence analysis
$\mathcal{W}(\hat{x}_{0:N}^{k+1},\hat{p}^{k+1}_{0:N})\leq\mathcal{W}~{}(\hat{x}_{0:N}^{k},\hat{p}^{k+1}_{0:N})\leq\mathcal{W}(\hat{x}_{0:N}^{k},\hat{p}^{k}_{0:N}).$ (31)
Therefore, the $0-1$ sequence $\hat{p}_{0:N}$ will iteratively minimize the objective function $\mathcal{W}$ in a monotonically nonincreasing way until $\hat{p}^{k+1}_{0:N}=\hat{p}^{k}_{0:N}$.
VII experiment
In this section, we introduce the comprehensive experimental results. First, we compare the update speeds of direct and iterative methods under different observation update rules. Then, We compare our GBSE with other detectors and estimators from the perspectives of detection rate of attacks and error in state estimation. Next, we demonstrate the detection capabilities of our GBSE algorithm in a multi-sensor system subjected to multiple attacks. The code is available at GitHub: https://github.com/RoboCubPilot/GB_secure_estimator.git.
Although our previous focus is on the single-sensor system, the derivation process and conclusions are also applicable to multi-sensor systems. For a multi-sensor system equipped with $M$ sensors, $y_{i,j}$ denotes the observation at sensor $j$ at time $i$, and the corresponding indicator is defined as $p_{i,j}$.
VII-A Comparison of direct approach and iterative approach
@@
First, in Fig. 7(a), we implement experiments to add observations along the timeline, which is the default observation update rule for most systems. The newly added observations include the observations from all sensors at time $k$ that $\mathcal{O}_{k}\setminus\mathcal{O}_{k-1}=\{k\}\otimes\{1,\cdots,M\}$. Additionally, in Fig. 7(b), we are interested in incorporating observations in the order of sensor indices. The rules for observation updates allow the system to update one after another based on the reliability of the sensors. The new observations add to the set from $k$-th sensor that $\mathcal{O}_{k}\setminus\mathcal{O}_{k-1}=\{0,\cdots,N\}\otimes\{k\}$. Moreover, in Fig. 7(c), we test randomly updated observation sets. In Algorithm 4, the update rule is unpredictable, simulating random observation masking. After masking part of the full set, we calculate the optimal estimate, then reintroduce masked observations, recording estimation times.
@@
The results show that the direct and iterative methods work best in different situations. The direct method is more efficient for timeline-based updates, while the iterative method is better when only a small, unordered part of the observation set is updated. The iterative method is also preferred when adding new sensor data to the observation set.
VII-B Compared with other detector and estimator
Consider the scale system with two sensors at interval $T=[0,N]$, where $x_{0}\sim\mathcal{N}(0,1),A=1,C_{1}=C_{2}=1,Q=0.5,R_{1}=R_{2}=2,N=20$. Let’s consider that sensor $1$ is functioning normally, while sensor $2$ has been injected with false information under attack at time $t$. Therefore,
@@ VII-B Compared with other detector and estimator
First, we implement the three common attack strategies in our experiment. Fig. 8(a) represents the application of random interference, where $e_{t}\sim N(0,\tilde{R})$. Fig. 8(b) indicates a constant effect on the sensor, where $e_{t}=\mu_{\text{constant}}$ starting from time $t=10$. Fig. 8(c) illustrates a uniformly increasing impact on the sensor, where $e_{t}=\frac{t}{N}\mu_{\text{max}}$.
@@
Then, we introduce the other detectors and estimators that our algorithm is compared against.
VII-B1 $\chi^{2}$ detector
The idea is to calculate the probability of the Kalman residuals and compare it with $\alpha_{\text{threshold}}$.
$\chi^{2}_{t}=\|y_{t}-C\hat{x}_{t|t-1}\|^{2}_{\sum_{r_{t}}^{-1}}\ \underset{}{\overset{}{\gtrless}}\ \alpha_{\text{threshold}}.$ (33)
VII-B2 CUSUM detector
It calculates the cumulative sum of the differences between the Kalman residuals and a reference $\delta_{\text{ref}}$ in different direction.
$S_{t}=\max(0,S_{t-1}+(r_{t}-\delta_{\text{ref}}))\ \underset{}{\overset{}{\gtrless}}\ \alpha_{\text{threshold}},$ (34)
where we set the $\delta_{\text{ref}}=0.5$.
VII-B3 Resilient estimator
The resilient estimator determines attacks by comparing $\chi^{2}$ with $\alpha_{\text{threshold}}$, but directly ignores those observations that exceed the $\alpha_{\text{threshold}}$.
VII-B4 Gaussian-Bernoulli Secure (GBS) estimator
Our proposed GBS estimator in Algorithm. 3 also needs to set the parameter $\alpha_{\text{threshold}}$ to determine low-probability events and the parameter $\tau_{\text{tolerant}}$ to tolerate their occasional occurrence.
We choose $\alpha_{\text{threshold}}=6$ and $\tau_{\text{tolerant}}=3$ that allows the detector and estimator to achieve optimal performance. We test their performance with varying attack intensities. The intensity is a topic of greater concern for us because, there is a trade-off between their stealthiness and the impact on state estimation. As shown in Fig. 9, we select six different levels of intensity for each type of attacks, ranging from mild to severe. For each attack intensity, we randomly simulate 10,000 observation sequences under attack.
@@ VII-B4 Gaussian-Bernoulli Secure (GBS) estimator
We further provide the detection success rate and estimation accuracy for different values of parameters $\alpha_{\text{threshold}}$ and $\tau_{\text{tolerant}}$ in Appendix B-B. The results show that the GBS estimator consistently outperforms other algorithms, regardless of the parameter choices.
VII-C Multiple attacks in multi-sensor system
In a multi-sensor system, as the number of sensors under attack increases, how to maintain the resilient state estimation and detect the attacked sensors is challenging. The reason is that, with the increase number of observations, attacks are more easily concealed among outliers, which are those occasional observations with significant errors.
@@ VII-C Multiple attacks in multi-sensor system
Suppose we deploy noise interference on sensors indexed from $1$ to $5$. When these attacks last from $i=0$ to $20$, the actual observation model becomes:
$y_{i,j}\sim\mathcal{N}(x_{i},\tilde{R}_{j}=100),j=1,\cdots,5.$
(a) system state and multiple sensor data
(b) map of observation noise
(c) map of estimated indicators
Figure 10: Characterization the measurement error by estimated observation indicators. (b) shows the true observation errors $\|y_{i,j}-C_{j}x_{i}\|^{2}_{R_{j}^{-1}}$, which are continuous values. Sensor $1$ to $5$, under attack, generated a large number of observations with significant errors. (c) shows our assessment of the reliability of each observation, which is boolean.
Comparing Fig. 10(b) and Fig. 10(c), their similarity illustrates that Gaussian-Bernoulli secure estimator effectively characterizes observation errors through indicators. Based on the estimated indicator sequence of each sensor, we can further distinguish between outliers and attacks.
VIII Conclusion
This paper has addressed the problem of sensor security by introducing a GBS estimator, which enables safety in terms of resilient estimation and attack detection. A dual-variable optimization problem, formulated as Problem 1, aims to estimate system states and sensor status using a Gaussian-Bernoulli observation model with indicators.
Starting from the Problem 2, we identify two practical state estimation problems underlying it: sequential state estimation with partial observations and estimation updates with disordered new observations. We then propose and derive direct and iterative state estimation methods based on the equivalent Problem 3 and Problem 4, which are suitable for different observation update patterns. Finally, we introduce the GBS estimator algorithm to solve Problem 1 and prove the existence and convergence of the solution.
In experiments, we compare the direct and iterative methods under different update patterns, simulate various sensor attacks, and show the GSB estimator’s superior detection and estimation. Lastly, we demonstrate that our Boolean indicator accurately matches actual observation errors.
Since the Gaussian-Bernoulli mixture model can effectively model systems that involve decision-making tasks, future work will focus on how to apply it to other control problems. One potential direction is to identify the system parameters while balancing the selection of effective data and the estimation of system parameters. These efforts will help us incorporate discrete objectives into the continuous system states.
Appendix A
A-A Proof of Lemma 1
@@
Observations contribute different weight to each state estimate, with the element of matrix $K_{\mathcal{O}}$ representing the optimal allocation of weights. For those observations $\{y_{t}|t\notin\mathcal{O}\}$, their weights in $K_{\mathcal{O}}$ equal to 00. As the set $\mathcal{O}$ changes, the corresponding optimal weights will also change.
A-B Proof of Theorem 2
@@ A-C Proof of Theorem 3
$\hat{x}_{k|\mathcal{O}^{*}}=\hat{x}_{k|\mathcal{O}}+P_{k|\mathcal{O}}C^{T}(CP_{k|\mathcal{O}}C^{T}+R)^{-1}(y_{k}-C\hat{x}_{k|\mathcal{O}})$ (60)
@@
Appendix B
B-A The limitation of $\chi^{2}$ based method
Consider a scalar system with single sensor at interval $T=[0,20]$, where $x_{0}\sim\mathcal{N}(0,1),A=1,C=1,Q=0.5,R=1.$ We simulate sensors in three different status, as shown in Fig. 11.
(a) three different sensor status
(b) state estimation from different sensors
(c) $\chi^{2}$ sequences from different sensors
Figure 11: Illustration of the limitation of $\chi^{2}$ based method. (a) displays three sets of sensor data: the blue represents normal observations. Compared to the blue observation, the green shows a deviation only at $t=10$, while the yellow deviates after $t=10$. (b)(c) shows their corresponding state estimates and $\chi^{2}_{t}$ values.
The experiment indicates that although the state estimates from Sensor 2 and Sensor 1 are very different, their $\chi^{2}$ curves are quite similar. This similarity arises because, after $t=10$, the state estimate from Sensor 2 becomes inaccurate, rendering the $\chi^{2}$ curve unreliable. Additionally, this phenomenon complicates the distinction between outliers and attacks, thus reducing the success rate of attack detection.
B-B Additional experiments with different algorithm parameters
Under different parameters, the algorithm’s estimation performance is roughly the same as shown in Fig. 9. The detection success rates corresponding to different parameters are illustrated in Fig. 12.
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=2$
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=4$
(a) random interference attack with different intensity
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=2$
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=4$
(b) constant bias attack with different intensity
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=2$
$\alpha_{\text{threshold}}=6,\tau_{\text{tolerant}}=4$
(c) uniformly increasing bias attack with different intensity

### v2

@@ 6.2 Gaussian-Bernoulli Secure (GBS) Estimator
$\arg\min_{\hat{p}_{0:N}}~\mathcal{W}=\sum_{i=0}^{N}[(1-\hat{p}_{i})\|y_{i}-C\hat{x}_{i}^{*}\|^{2}_{R^{-1}}+\hat{p}_{i}\alpha]$ (19)
@@
We can update $\hat{p}_{k}$ by solving the optimization problem:
@@
As described in Algorithm 3, we iteratively calculate equation (18) and (19), and improve the optimality through equation (23).
@@
1:system parameters $\{A,C,Q,R,\alpha,\tau\}$
2:initial state $(\bar{x}_{0},\bar{P}_{0})$ and observations $y_{0:N}$
[…]
4:Initialize $\hat{x}_{0:N}^{0},\hat{p}_{0:N}^{0}$ using Algorithm 2
5:Set count $k=0$ and ${\mathcal{O}^{0}}=\{i\mid\hat{p}_{i}^{0}=0\}$
6:repeat
7:  $k\leftarrow k+1$
8:  for $i=0$ to NN do
9:   $\hat{p}_{i}^{k}=\textbf{bool}(\|y_{i}-C\hat{x}_{i}^{*}\|^{2}_{R^{-1}}>\alpha)$
[…]
11:  Update $\mathcal{O}^{k}=\{i\mid\hat{p}_{i}^{k}=0\}$
12:  if $\mathcal{O}^{k}=\mathcal{O}^{k-1}$ then
13:   for $i=0$ to NN do
14:     $\hat{p}_{i}^{*}=\textbf{bool}((\min\mathcal{L}_{\mathcal{O}_{i}^{+}}-\min\mathcal{L}_{\mathcal{O}_{i}^{-}})>\alpha)$
15:     if $\hat{p}_{i}^{*}\neq\hat{p}_{i}^{k}$ then
16:      $\hat{p}_{i}^{k}\leftarrow\neg\hat{p}_{i}^{k}$, update $\mathcal{O}^{k}$
17:      break
18:     end if
19:   end for
[…]
21:  if $\mathcal{O}^{k}=\mathcal{O}^{k-1}$ then
22:   break
23:  else
24:   Update optimal estimates $\hat{x}_{0:N}$
[…]
26:until
27:if $(N-|\mathcal{O}^{k}|)>\tau$ then
28:  Print: info("attack alarm in this sensor!")
29:end if
6.3 Convergence analysis
In this subsection, we prove the convergence of our proposed GBS estimator algorithm.
First, we demonstrate the existence of the solution to our target MIP problem presented in (7). As a 010-1 sequence $\hat{p}_{0:N}\in\{0,1\}^{(N+1)}$, there are a total of $2^{(N+1)}$ cases. For $\{0,1\}^{(N+1)}\leftrightarrow 2^{\{0,\cdots,N\}}$, we can define a mapping from the indicator set to the index set that
@@ 6.3 Convergence analysis
$\mathcal{W}(\hat{x}_{0:N}^{k+1},\hat{p}^{k+1}_{0:N})\leq\mathcal{W}~(\hat{x}_{0:N}^{k},\hat{p}^{k+1}_{0:N})\leq\mathcal{W}(\hat{x}_{0:N}^{k},\hat{p}^{k}_{0:N}).$
Therefore, the 010-1 sequence $\hat{p}_{0:N}$ will iteratively minimize the objective function $\mathcal{W}$ in a monotonically nonincreasing way until $\hat{p}^{k+1}_{0:N}=\hat{p}^{k}_{0:N}$.
7 Numerical Simulations
In this section, we introduce the comprehensive experimental results. First, we compare the update speeds of direct and iterative methods under different observation update rules. Then, we compare our GBSE with other detectors and estimators from the perspectives of detection and estimation performance. Next, we demonstrate the detection capabilities of our GBSE algorithm in a multi-sensor system.
Although our previous focus is on the single-sensor system, the derivation process and conclusions are also applicable to multi-sensor systems. For a multi-sensor system equipped with $M$ sensors, $y_{i,j}$ denotes the observation at the sensor $j$ at time $i$, and the corresponding indicator is defined as $p_{i,j}$.
7.1 Comparison of Direct and Iterative Approach
@@
First, in Fig. 7(a), we implement experiments to add observations along the timeline, which is the default observation update rule for most systems. The newly added observations include the observations from all sensors at time $k$ that $\mathcal{O}_{k}\setminus\mathcal{O}_{k-1}=\{k\}\otimes\{1,\cdots,M\}$. Additionally, in Fig. 7(b), we are interested in incorporating observations in the order of sensor indices. The rules for observation updates allow the system to update one after another based on the reliability of the sensors. The new observations add to the set from $k$-th sensor that $\mathcal{O}_{k}\setminus\mathcal{O}_{k-1}=\{0,\cdots,N\}\otimes\{k\}$. Moreover, in Fig. 7(c), we test randomly updated observation sets, similar to the setup in Algorithm3. We first hide part of the observations and compute the optimal estimate. Then, we add the masked observations back and record the computation time required to update the estimate.
@@
The results show that the direct and iterative methods work best in different situations. The direct method is more efficient for timeline-based updates, while the iterative method is better when a small, unordered part of the observation set is updated. The iterative method is also preferred when adding new sensors.
7.2 Compared with Other Detectors and Estimators
Consider the scale system with two sensors at interval $T=[0,N]$, where $x_{0}\sim\mathcal{N}(0,1),A=1,C_{1}=C_{2}=1,Q=0.5,R_{1}=R_{2}=2,N=20$. Let’s consider that sensor $1$ is functioning normally, while sensor $2$ has been injected with false information under attack at time $t$.
@@ 7.2 Compared with Other Detectors and Estimators
[…]
@@
The following detectors and estimators are considered for comparison:
  • $\chi^{2}$ detector calculates the probability of the Kalman residuals and compares it with a threshold $\alpha_{\text{threshold}}$:
    $\chi^{2}_{t}=\|y_{t}-C\hat{x}_{t|t-1}\|^{2}_{\Sigma_{r_{t}}^{-1}}\ \underset{}{\overset{}{\gtrless}}\ \alpha_{\text{threshold}}.$ (29)
  • CUSUM detector computes the cumulative sum of the deviations of the Kalman residuals from a reference value $\delta_{\text{ref}}=0.5$ in a given direction:
    $S_{t}=\max(0,S_{t-1}+(r_{t}-\delta_{\text{ref}}))\ \underset{}{\overset{}{\gtrless}}\ \alpha_{\text{threshold}}.$ (30)
  • Resilient estimator detects attacks by checking whether the $\chi^{2}$ value exceeds $\alpha_{\text{threshold}}$, and ignores the corresponding observations if so.
  • Gaussian-Bernoulli Secure (GBS) estimator sets the threshold $\alpha_{\text{threshold}}$ to detect low-probability events and introduces a tolerance parameter $\tau_{\text{tolerant}}$ to allow for occasional anomalies.
We choose $\alpha_{\text{threshold}}=6$ and $\tau_{\text{tolerant}}=3$ that allows each detector and estimator to achieve optimal performance. We evaluate their performance under varying attack intensities. Intensity is of particular interest to us, as there is a trade-off between the stealthiness of the attack and its impact on state estimation. As shown in Fig. 9, we select six levels of intensity for each type of attack, ranging from mild to severe. For each level, 10,000 observation sequences under attack are randomly generated.
@@
7.3 Multiple Attacks in Multi-sensor System
In multi-sensor systems, detecting attacks and ensuring resilient estimation becomes harder as more sensors are compromised. The difficulty arises from the fact that, as the volume of observations grows, malicious data can be more easily hidden among outliers—rare but significant deviations.
@@ 7.3 Multiple Attacks in Multi-sensor System
Suppose we deploy noise interference on sensors indexed from $j^{*}=1$ to $5$. When these attacks last from $i=0$ to $20$, the actual observation model becomes:
$y_{i,j^{*}}\sim\mathcal{N}(x_{i},\tilde{R}_{j^{*}}=100),j^{*}=1,\cdots,5.$
Comparing Fig. 10(a) and Fig. 10(b), their similarity illustrates that Gaussian-Bernoulli secure estimator effectively characterizes observation errors via indicators. Based on the estimated indicator sequence of each sensor, we can further distinguish attacks from outliers.
(a) observation noise map
(b) estimated indicators map
Figure 10: Characterization the measurement error by estimated observation indicators. (a) shows the true observation errors $\|y_{i,j}-C_{j}x_{i}\|^{2}_{R_{j}^{-1}}$, which are continuous values. Sensor 11 to $5$, under attack, generated several observations with significant errors. (b) shows our assessment of the reliability of each observation, which is boolean.
8 Conclusion
We revisit the sensor attack problem from the perspective of system-environment interaction. Unlike conventional systems, interactive systems emphasize that observations not only contain information strongly correlated with the system’s internal state, but also carry information determined by the external environment. Under potential sensor attacks, an observation variable is determined by the system state during normal operation, but by the external environment when the sensor is compromised. Therefore, in interactive systems, state estimation can be viewed as the process of decoding system information, while attack detection becomes the task of decoding environmental information.
There are two main challenges in decoding such observation information. The first challenge lies in how to effectively characterize the environmental attack information, especially when the attack model is unknown, to facilitate subsequent decoding. The second challenge is that information from these two distinct sources is often entangled in a single observation variable, making it necessary to decouple them before decoding.
The central motivation and contribution of our work is to address these two challenges. First, we propose the Bernoulli-Gaussian observation model assumption, where the system’s internal information is modeled using a Gaussian distribution, and unknown environmental attacks are modeled using a Bernoulli distribution. Based on this modeling assumption, we formulate a dual-variable optimization problem to jointly infer the system state and the observation indicators. Finally, we propose the GBS estimator, an efficient algorithm capable of decoupling and decoding both the system state and the attack information from the observations.
Appendix A APPENDIX
A.1 Proof of Lemma 1
@@
Observations contribute different weight to each state estimate, with the element of matrix $K_{\mathcal{O}}$ representing the optimal allocation of weights. For those observations $\{y_{t}|t\notin\mathcal{O}\}$, their weights in $K_{\mathcal{O}}$ equal to 0. As the set $\mathcal{O}$ changes, the corresponding optimal weights will also change.
A.2 Proof of Theorem 2
@@ A.3 Proof of Theorem 3
$\hat{x}_{k|\mathcal{O}^{*}}=\hat{x}_{k|\mathcal{O}}+P_{k|\mathcal{O}}C^{T}(CP_{k|\mathcal{O}}C^{T}+R)^{-1}(y_{k}-C\hat{x}_{k|\mathcal{O}})$